# Connected Network Model for the Mechanical Loss of Amorphous Materials

### v2

@@ Abstract
Dissipation in amorphous solids at low frequencies is commonly attributed to activated transitions of isolated two-level systems (TLS) that come in resonance with elastic or electric fields. Materials with low mechanical or dielectric loss are urgently needed for applications in gravitational wave detection, high precision sensors, and quantum computing. Using atomistic modeling, we explore the energy landscape of amorphous silicon and titanium dioxide, and find that the pairs of energy minima that constitute single TLS form a sparsely connected network with complex topologies. Motivated by this observation, we develop an analytically tractable theory for mechanical loss of the full network from a nonequilibrium thermodynamic perspective. We demonstrate that the connectivity of the network introduces new mechanisms that can both reduce low frequency dissipation through additional low energy relaxation pathways, and increase dissipation through a broad distribution of energy minima. As a result, the connected network model predicts mechanical loss with distinct frequency profiles compared to the isolated TLS model. This not only calls into question the validity of the TLS model, but also gives us many new avenues and properties to analyze for the targeted design of low mechanical loss materials.
@@ I Introduction
The physical origin of energy dissipation in amorphous films is ascribed to the presence of two-level systems (TLS), whose thermally activated transitions come in resonance with a mechanical wave [5]. The TLS represent atomic level transitions within amorphous solids; for example, amorphous silicon (a-Si) may undergo bond hopping transitions [6]. The two distinct atomic configurations may be associated with pairs of stable minima (inherent structures) in an energy landscape connected by a transition path. External fields drive the TLS out of equilibrium, resulting in internal friction upon relaxation [7, 5, 8]. Mechanical loss is caused by elastic deformation of the atoms constituting the TLS, while electromagnetic waves coupling to the TLS’ dipole moment cause dielectric loss.
@@
We explore the consequences of an ensemble of TLS forming connected networks for the mechanical loss of amorphous solids. Our focus lies on classical thermal transitions as they relate to the mechanical loss near room temperature, and we take a non-equilibrium thermodynamic perspective [10]: modeling the heat dissipated from a driven process using master equation dynamics [11] of the discrete state network. In the limit of linear response, we arrive at a rigorous generalization of the 50 year old single TLS model that accounts for the additional relaxation paths afforded by the connectivity of the network. Through simple model systems and full atomistic modeling, we reveal mechanisms by which such an ensemble of connected TLS can significantly reduce mechanical loss in the frequency band of $10^{1}-10^{4}\,$z where ground-based gravitational wave detection takes place.
Figure 1: Left: First $100$ nodes of the connected network of inherent structures observed in samples of amorphous silicon (a-Si) and amorphous titanium dioxide (a-TiO2). The color of the dots and lines correspond to the energy of the minima $E_{i}$ and the average barrier height $V_{ij}-(E_{i}+E_{j})/2$ respectively. Middle: Distributions of energy minima $E_{i}$ (black solid curve), energy asymmetries $\Delta_{ij}=E_{j}-E_{i}$ (red dotted), and barrier heights $V_{ij}$ (blue dashed) between inherent structures ($i$,$j$) for a-Si (top) and a-TiO2 (bottom). The energetics for a pair of connected inherent structures is shown schematically in the inset. Right: Fraction of the number of basis cycles and probability distribution of the degree (number of connections) of nodes in the network with the dashed line a linear least squares fit to the a-Si data with a slope (scaling exponent) of $-2.84\pm 0.04$. Error bars in the degree distribution are the standard error of the mean from logarithmically binned averages. Basis cycles are the minimal set of closed loops in the network, with examples of a three-state cycle in a-TiO2 and a four-state cycle in a-Si shown in the inset.
@@ II Energy landscapes from atomistic simulations
We begin by exploring the energy landscapes of in-silico samples of a-Si and a-TiO2 as described in section ’Materials and Methods’. Through thermal molecular dynamics search trajectories, we find inherent structures (stable minima of the energy landscape) connected by thermodynamic pathways determined through nudged elastic band (NEB) calculations. Representing energy minima as nodes and thermodynamic transitions as edges, the energy landscapes form sparsely connected networks (Fig. 1, left panel). Each connected pair of nodes would be considered an individual two-level system within the TLS model (Fig. 1, middle panel inset). Although only the first 100 nodes are shown for clarity, the observed networks often contain thousands of inherent structures (nodes). In a-Si, the percentage of the network that is connected is $\sim 0.05\%$, while in a-TiO2 the connectivity of the largest networks is $\sim 0.14\%$.
The distributions of both the energy minima and the energy barriers that link them are broad (middle panel), reflecting the structural disorder. Locally, we find that nearby energy minima often from closed loops or cycles that can be viewed as multiple TLS linked together (right panel). The degree distribution of these networks follows a power law, indicating that the networks are scale free. Similar scale-free networks are found in Lennard-Jones clusters [12], while the energy landscape picture has been invoked to explain numerous phenomena associated with the glass transition  [9, 13, 14].
@@ III.1 Thermodynamics
We separate the total energy of the system into the elastic energy of the acoustic wave of frequency $\omega$ traveling through the material and the internal energy of the connected network of inherent structure as $\mathcal{U}_{\rm tot}=\mathcal{U}_{\rm elastic}+\mathcal{U}_{\rm CN}$. Assuming an isotropic and linear elastic material with oscillating strain $\epsilon(t)=\epsilon_{0}\sin(\omega t)$, the elastic energy is
$\mathcal{U}_{\rm elastic}(t)=\mathcal{V}C\epsilon_{0}^{2}\sin^{2}(\omega t)\ ,$ (2)
@@
where $\boldsymbol{1}$ is a vector of ones and throughout a bold symbol represent a vector with elements spanning the inherent structures; i.e. $E_{i}$ is the energy of inherent structure $i$. The coupling between the inherent structure and applied strain is separated into a constant term $\alpha$ and an inherent structure dependent term $\boldsymbol{\Gamma}$ with amplitude $\gamma_{0}$. The constant term will not affect the dynamics since it will affect all energies equally. The average energy of the network is the number of inherent structures $N$ times the average energy per inherent structure
@@
$\dot{\mathcal{W}}=\dot{\mathcal{U}}_{\rm elastic}+N\boldsymbol{P}_{t}\cdot\dot{\boldsymbol{E}}(t)\ ,$ (6)
@@ III.3 Linear Response (Weak Limit)
Although the equations of the previous section (Eq. (7) and Eq. (8)) are sufficient to determine the energy dissipation and hence the mechanical loss of the system; considerable simplification can be made if we assume the amplitude of oscillations are small. This is consistent with the small amplitude assumption made in the TLS model [5] and should be sufficient for applications to GWD where the magnitude of vibrations in the system are extremely small [2].
@@
$\boldsymbol{P}^{(1)}_{t}$ $=-\frac{\beta\epsilon_{0}\gamma_{0}}{2}\left[\boldsymbol{A}\sin(\omega t)+\boldsymbol{B}\cos(\omega t)\right]$ (17)
@@
The relaxation times $\tau_{j}$ are related to the eigenvalues $\lambda_{j}$ of $R^{(0)}$ as $\tau_{j}\equiv\lambda_{j}^{-1}$ with corresponding eigenvectors forming the columns of the eigenvector matrix $M$.
@@
The TLS model corresponds to a network where every pair of inherent structures are isolated. For this network there is no global equilibrium and each pair of connected inherent structures must be treated independently. In Supplemental Material II, we calculate the quality factor for a network of two inherent structures, which when summed over all independent pairs yields a quality factor consistent with the popular TLS model [5]
$Q_{\rm TLS}^{-1}=\frac{\beta}{4\mathcal{V}C}\sum_{i}\gamma_{i}^{2}\operatorname{sech}^{2}\left(\frac{\beta\Delta_{i}}{2}\right)\frac{\omega\tau_{i}}{1+\omega^{2}\tau_{i}^{2}}\ ,$ (20)
with $\Delta_{i}$ the energy difference between the two inherent structures in TLS $i$, $\gamma_{i}=\gamma_{0}\Gamma_{i}$ the TLS deformation potential, and TLS relaxation time
@@ IV.1 Cycles and Connectivity: Four-State Networks
Figure 2: Scaled mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ as a function of frequency for a) a four-state chain and b) a four-state cycle. Two different curves are shown: the TLS approximation (20) (black solid line), and the connected network  (19) (red dashed). In all cases the large energy barrier is $0.8$eV with the remaining barriers $0.6$eV.
Our atomistic samples of amorphous materials exhibit a significant fraction of basis cycles (Fig. 1), with four-state the most frequent. We therefore consider a minimal four-state network that consists of discrete states $i$ with energy minima $E_{i}$ and barriers $V_{ij}$ as depicted in Fig. 2. All states separated by a barrier are connected with transitions and dynamics given by eq. (8). We estimate mechanical loss from eq. (19) and eq. (20) and in all cases the temperature is $300K$.
@@
$R_{0}^{\rm chain}=k_{0}\begin{bmatrix}-x&x&0\\ x&-(x+1)&1\\ 0&1&-1\end{bmatrix}\ ,$ (22)
@@
$R_{0}^{\rm cycle}=k_{0}\begin{bmatrix}-(x+1)&x&1\\ x&-(x+1)&1\\ 1&1&-2\end{bmatrix}\ .$ (23)
In both cases the zero eigenvalue corresponds to the equilibrium distribution with equal probability in each state. For the chain, the remaining two eigenvalues are $\lambda_{2}^{\rm chain}\approx-xk_{0}$ and $\lambda_{3}^{\rm chain}\approx-2k_{0}$ with eigenvectors $[-2,1,1]$ and $[0,-1,1]$ respectively. The former results in a low-frequency peak in dissipation since $x$ is assumed to be small. For the cycle, the remaining two eigenvalues are $\lambda_{2}^{\rm cycle}\approx-3k_{0}$ and $\lambda_{3}^{\rm cycle}\approx-k_{0}$ with eigenvectors $[-1/2,-1/2,1]$ and $[-1,1,0]$ respectively. For the cycle the eigenvalues are approximately independent of the small relative transition rate $x$, effectively removing the low frequency peak observed in the chain. Similar calculations for a four-state chain yields eigenvalues $0,-4/3xk_{0},-k_{0},-3k_{0}$ and for a four-state cycle $20,-2,-\sqrt{2}-2,\sqrt{2}-$ 2, showing an identical phenomenon.
Performing the same comparison between the four-state chain and cycle at fixed frequency and variable temperature would show the same qualitative trends. The low frequency peak becomes a high temperature peak in the mechanical loss-Temperature plane and the high frequency peak becomes a low temperature peak. The overall effect of connecting four states in a cycle can therefore be equally thought of as suppressing high temperature relaxation modes at fixed frequency.
@@ IV.3 Energy Distribution: Four-State Networks
Figure 4: Mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ as a function of frequency. Two different colors (shades) of curves are shown: the TLS approximation (20) (black), and the connected network linear-response approximation (19) (red). The energy of the center two minima is $0,0.15,0.3$ eV for solid, dashed, dotted, and dot-dashed curves respectively. In all cases the energy barriers are $0.4$eV above the lower energy level connecting two states.
In the previous sections, we showed how the connectivity of the network can completely eliminate low frequency modes; however, if the distribution of energy minima is broad compared to the distribution of energy barriers the opposite effect occurs—the connected network results in low frequency modes not observed in the independent TLSs. As before, if the connectivity is increased far enough the low frequency modes are once again eliminated.
To illustrate this phenomena we consider a four-state network with a “mountain” landscape as depicted in Fig. 4. Since the TLS model only compares pairs of energies, it has no sense of a global energy scale. Therefore, increasing the energy of the two minima in the center simply increases the asymmetry of two of the three TLS, decreasing dissipation without altering the frequency spectrum.
@@
$R_{0}^{\rm mountain}=k_{0}\begin{bmatrix}-1&x&0&0\\ 1&-(x+1)&1&0\\ 0&1&-(x+1)&1\\ 0&0&x&-1\end{bmatrix}\ ,$ (25)

### v3

@@ Abstract
Dissipation in amorphous solids at low frequencies is commonly attributed to activated transitions of isolated two-level systems (TLS) that come in resonance with elastic or electric fields. Materials with low mechanical or dielectric loss are urgently needed for applications in gravitational wave detection, high precision sensors, and quantum computing. Using atomistic modeling of amorphous silicon and titanium dioxide, we find that their energy landscape is better represented by a connected network of inherent structures than a collection of isolated TLS. Each connection is a single energy barrier between two minima, and a network is connected if all states (minima) can be reached from any given state. Motivated by this observation, we develop an analytically tractable theory for mechanical loss of the full network from a nonequilibrium thermodynamic perspective. We demonstrate that the connectivity of the network introduces new mechanisms that can both reduce low frequency dissipation through additional low energy relaxation pathways, and increase dissipation through a broad distribution of energy minima. As a result, the connected network model predicts mechanical loss with distinct frequency profiles compared to the isolated TLS model. This not only calls into question the validity of the TLS model, but also gives us many new avenues and properties to analyze for the targeted design of low mechanical loss materials.
@@ I Introduction
The physical origin of energy dissipation in amorphous films is ascribed to the presence of two-level systems (TLS), whose thermally activated transitions come in resonance with a mechanical or electromagnetic wave [8]. The TLS represent atomic level transitions within amorphous solids; for example, amorphous silicon (a-Si) may undergo bond hopping transitions [21]. The two distinct atomic configurations may be associated with pairs of stable minima (inherent structures) in an energy landscape connected by a transition path. External fields drive the TLS out of equilibrium, resulting in internal friction upon relaxation [32, 8, 6]. Mechanical and dielectric loss are caused by small displacements of atoms constituting the TLS, with mechanical waves coupling to the TLS in mechanical loss and electromagnetic waves coupling to the TLS’ dipole moment in dielectric loss.
@@
We explore the consequences of accounting for the rugged and high-dimensional energy landscape through a connected network model. A connected network is a discrete-state representation of the high dimensional energy landscape, with each node representing an energy minimum and each connection (edge) a single energy barrier. In a connected network, every state is accessible from any given state, giving a consistent thermodynamic description of the energy landscape. Our focus lies on classical thermal transitions as they relate to the mechanical loss near room temperature, and we take a non-equilibrium thermodynamic perspective [36]: modeling the heat dissipated from a driven process using master equation dynamics [14] of the discrete state network. In the limit of linear response, we arrive at a rigorous generalization of the 50 year old single TLS model that accounts for the additional relaxation paths afforded by the connectivity of the network. Through simple model systems and full atomistic modeling, we reveal mechanisms by which activated transitions in a larger connected network of energy minima can significantly reduce mechanical loss in the frequency band of $10^{1}-10^{4}\,$Hz where ground-based gravitational wave detection takes place.
Figure 1: Left: First $100$ nodes of the connected network of inherent structures observed in samples of amorphous silicon (a-Si) and amorphous titanium dioxide (a-TiO2). The color of the dots and lines correspond to the energy of the minima $E_{i}$ and the average barrier height $V_{ij}-(E_{i}+E_{j})/2$ respectively. Middle: Distributions of energy minima $E_{i}$ (black solid curve), magnitude of energy asymmetries $\Delta_{ij}=E_{j}-E_{i}$ (red dotted), and barrier heights $V_{ij}$ (blue dashed) between inherent structures (ii,$j$) for a-Si (top) and a-TiO2 (bottom). The energetics for a pair of connected inherent structures is shown schematically in the inset. Right: Fraction of the number of basis cycles and probability distribution of the degree (number of connections) of nodes in the network with the dashed line a linear least squares fit to the a-Si data with a slope (scaling exponent) of $2.84\pm 0.04$. Error bars in the degree distribution are the standard error of the mean from logarithmically binned averages. Basis cycles are the minimal set of closed loops in the network, with examples of a three-state cycle in a-TiO2 and a four-state cycle in a-Si shown in the inset.
@@ II Energy landscapes from atomistic simulations
We begin by exploring the energy landscapes of in-silico samples of a-Si and a-TiO2 as described in section Materials and Methods. Through thermal molecular dynamics search trajectories, we find inherent structures (stable minima of the energy landscape) connected by thermodynamic pathways determined through nudged elastic band (NEB) calculations. Representing energy minima as nodes and thermodynamic transitions as edges, the energy landscapes form sparsely connected networks (Fig. 1, left panel). Each connected pair of nodes would be considered an individual two-level system within the TLS model (Fig. 1, middle panel inset). Although only the first 100 nodes are shown for clarity, the observed networks often contain thousands of inherent structures (nodes). In a-Si, the percentage of the network that is connected is estimated to be $\sim 0.05\%$, while in a-TiO2 the connectivity of the largest networks is $\sim 0.14\%$. Although the energy landscape search via a random walk is not guaranteed to explore all edges or inherent structures, we find these connectivity values robust and converged within our numerical protocol (see Supplemental Material IV).
The distributions of both the energy minima and the energy barriers that link them are broad (middle panel), reflecting the structural disorder. Locally, we find that nearby energy minima often form closed loops or cycles (right panel). The degree distribution of these networks follows a power law, indicating that the networks are scale free. Similar scale-free networks are found in Lennard-Jones clusters [11], while the energy landscape picture has been invoked to explain numerous phenomena associated with the glass transition  [35, 1, 4].
Traditionally, each transition between inherent structures is treated as a new independent TLS [9]. Since the inherent structure describes a minimal energy configuration of the entire molecular system, in principle the configurations of $N$ TLS can result in $2^{N}$ inherent structures. The network description of the TLS model with $N$ pairs of states would form a $2^{N}$ state hypercubic network with delta function degree distribution $\delta({\rm degree}-N)$ and only 4-state basis cycles. The mechanical loss of such a network is identical to the TLS model only if the transitions and their response to mechanical perturbations are independent from each other. For example, two spatially separated bond-hop transitions that can oscillate independently would form a $2^{2}=4$ state network with one four-state cycle (Supplemental Material II).
@@ III.1 Thermodynamics
We separate the total energy of the system into the elastic energy of the acoustic wave of frequency $\omega$ traveling through the material and the internal energy of the connected network of inherent structure as $\mathcal{U}_{\rm tot}=\mathcal{U}_{\rm elastic}+\mathcal{U}_{\rm CN}$. Assuming an isotropic and linear elastic material with oscillating strain $\epsilon(t)=\epsilon_{0}\sin(\omega t)$, the elastic work is
$\mathcal{W}_{\rm elastic}(t)=\mathcal{V}C\epsilon_{0}^{2}\sin^{2}(\omega t)\ ,$ (2)
@@
where $\boldsymbol{1}$ is a vector of ones and throughout a bold symbol represent a vector with elements spanning the inherent structures; i.e. $E_{i}$ is the energy of inherent structure $i$. The coupling between the inherent structure and applied strain is separated into a constant term $\alpha$ and an inherent structure dependent term $\epsilon_{0}\gamma_{0}\boldsymbol{\Gamma}/2$, with strain amplitude $\epsilon_{0}\gamma_{0}$ and dimensionless inherent structure coupling $\boldsymbol{\Gamma}$. The constant term will not affect the dynamics since it will affect all energies equally. The average energy of the network is the number of inherent structures $N$ times the average energy per inherent structure
@@
$\dot{\mathcal{W}}=\dot{\mathcal{W}}_{\rm elastic}+N\boldsymbol{P}_{t}\cdot\dot{\boldsymbol{E}}(t)\ ,$ (6)
@@ III.3 Linear Response (Weak Limit)
Although the equations of the previous section (Eq. (7) and Eq. (8)) are sufficient to determine the energy dissipation and hence the mechanical loss of the system, considerable simplification can be made if we assume the amplitude of oscillations are small. This is consistent with the small amplitude assumption made in the TLS model [8] and should be sufficient for applications to GWD where the magnitude of vibrations in the system are extremely small [37].
@@
$\boldsymbol{P}^{{\rm weak}}_{t}$ $=-\frac{\beta\epsilon_{0}\gamma_{0}}{2}\left[\boldsymbol{A}\sin(\omega t)+\boldsymbol{B}\cos(\omega t)\right]$ (17)
@@
The relaxation times $\tau_{j}$ are related to the eigenvalues $\lambda_{j}$ of $R_{0}$ as $\tau_{j}\equiv\lambda_{j}^{-1}$ with corresponding eigenvectors forming the columns of the eigenvector matrix $M$.
@@
The TLS model corresponds to a network where every pair of inherent structures are isolated. For this network there is no global equilibrium and each pair of connected inherent structures must be treated independently. In Supplemental Material II, we calculate the quality factor for a network of two inherent structures, which when summed over all independent pairs yields a quality factor consistent with the popular TLS model that sums over individual TLS contributions  [8]
$Q_{\rm TLS}^{-1}=\frac{\beta\gamma_{0}^{{2}}}{4\mathcal{V}C}\sum_{i}\Gamma_{i}^{2}\operatorname{sech}^{2}\left(\frac{\beta\Delta_{i}}{2}\right)\frac{\omega\tau_{i}}{1+\omega^{2}\tau_{i}^{2}}\ ,$ (20)
with $\Delta_{i}$ the energy difference between the two inherent structures in TLS $i$ and $\Gamma_{i}$ their deformation potential, as well as the TLS relaxation time
@@
In going from Eq. 19 to Eq. 20 the $\operatorname{sech}^{2}$ term arises from the eigenvectors $M$ and the relative occupation probability of the two TLS through the Boltzmann factors in $P^{\rm eq}$.
The TLS mechanical loss can also be derived directly from the network description. Assuming every TLS $i$ has equal occupation probability and transitions independently, their transition rates $R^{\rm TLS}_{i}$ form $2\times 2$ matrices. The global transition rate matrix for $N$ independent Markov jump processes is the Kronecker sum of the constituent rate matrices $R=\oplus_{i=1}^{N}R^{\rm TLS}_{i}$. The eigenvectors of the $2^{N}$ element matrix $R$ are the Kronecker product of the individual TLS eigenvectors with eigenvalues the corresponding sum of the individual TLS eigenvalues [14].
This can give rise to correlated eigenmodes involving multiple TLS eigenmodes. If each TLS also responds independently to mechanical perturbations, then the global coupling terms can be expressed as the sum of the individual TLS couplings. The sum of one-body functions never produces two-body (or higher) tensor components, and only the individual TLS eigenmodes contribute to mechanical loss. The special case of $N=2$ is treated in Supplemental Material II.
There are three assumptions that must be true for the TLS model to be valid: i) the equilibrium distribution $P^{\rm eq}$ is constant across different TLS, ii) the TLSs transition independently, and iii) the TLS couple independently to mechanical perturbations. These are strong assumptions given the disordered structure of amorphous solids and the potential for long range elastic couplings. We find that relaxing these assumptions dramatically changes the predicted mechanical loss.
@@ IV.1 Cycles and Connectivity: Four-State Networks
Figure 2: Scaled mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ as a function of frequency for a) a four-state chain and b) a four-state cycle. Two different curves are shown: the TLS approximation (20) (black solid line), and the connected network  (19) (red dashed). In all cases the large energy barrier is 0.80.8eV with the remaining barriers 0.60.6eV.
Our atomistic samples of amorphous materials exhibit a significant fraction of basis cycles (Fig. 1), with four-state cycles the most frequent. We therefore consider a minimal four-state network that consists of discrete states $i$ with energy minima $E_{i}$ and barriers $V_{ij}$ as depicted in Fig. 2. All states separated by a barrier are connected with transitions and dynamics given by eq. (8). We estimate mechanical loss from eq. (19) and eq. (20) and in all cases the temperature is $300K$.
@@
$R_{0}^{\rm chain}=k_{0}\begin{bmatrix}-x&x&0\\ x&-(x+1)&1\\ 0&1&-1\end{bmatrix}\$, (22)
@@
$R_{0}^{\rm cycle}=k_{0}\begin{bmatrix}-(x+1)&x&1\\ x&-(x+1)&1\\ 1&1&-2\end{bmatrix}\$. (23)
In both cases the zero eigenvalue corresponds to the equilibrium distribution with equal probability in each state. For the chain, the remaining two eigenvalues are $\lambda_{2}^{\rm chain}\approx-xk_{0}$ and $\lambda_{3}^{\rm chain}\approx-2k_{0}$ with eigenvectors [2,1,1][-2,1,1] and [0,1,1][0,-1,1] respectively. The former results in a low-frequency peak in dissipation since $x$ is assumed to be small. For the cycle, the remaining two eigenvalues are $\lambda_{2}^{\rm cycle}\approx-3k_{0}$ and $\lambda_{3}^{\rm cycle}\approx-k_{0}$ with eigenvectors [1/2,1/2,1][-1/2,-1/2,1] and [1,1,0][-1,1,0] respectively. For the cycle the eigenvalues are approximately independent of the small relative transition rate $x$, effectively removing the low frequency peak observed in the chain. Similar calculations for a four-state chain yields eigenvalues $0,-4/3xk_{0},-k_{0},-3k_{0}$ and for a four-state cycle $0,-2,-\sqrt{2}-2,\sqrt{2}-2$, showing an identical phenomenon.
Performing the same comparison between the four-state chain and cycle at fixed frequency and variable temperature would show the same qualitative trends. The low frequency peak becomes a high temperature peak in the mechanical loss-Temperature plane and the high frequency peak becomes a low temperature peak. The overall effect of connecting four states in a cycle can therefore be equally thought of as suppressing high temperature relaxation modes at fixed frequency. The next simplest structure formed by a four-state network would allow for connections between any of the four nodes. This results in a combination of three- and four-state cycles that behave similarly to the four-state cycle tested in this section.
@@ IV.3 Energy Distribution: Four-State Networks
Figure 4: Mechanical loss $Q^{-1}\mathcal{V}C/(\beta\gamma_{0}^{2})$ as a function of frequency. Two different colors (shades) of curves are shown: the TLS approximation (20) (black), and the connected network linear-response approximation (19) (red). The energy of the center two minima is $0,0.15,0.3$ eV for solid, dashed, and dotted curves respectively. In all cases the energy barriers are 0.40.4eV above the lower energy level connecting two states.
In the previous section, we showed how the connectivity of the network can completely eliminate low frequency modes; however, if the distribution of energy minima is broad compared to the distribution of energy barriers the opposite effect occurs—the connected network results in low frequency modes not observed in the independent TLSs. As before, if the connectivity is increased far enough the low frequency modes are once again eliminated.
To illustrate this phenomenon we consider a four-state network with a “mountain” landscape as depicted in Fig. 4. Since the TLS model only compares pairs of energies, it has no sense of a global energy scale. Therefore, increasing the energy of the two minima in the center simply increases the asymmetry of two of the three TLS, decreasing dissipation without altering the frequency spectrum.
@@
$R_{0}^{\rm mountain}=k_{0}\begin{bmatrix}-1&x&0&0\\ 1&-(x+1)&1&0\\ 0&1&-(x+1)&1\\ 0&0&x&-1\end{bmatrix}\$, (25)